# A $2$-distance $(2\Delta+7)$-coloring of planar graphs

### v2

@@ 2.3. The case $\Delta=8$
Let $v$ be a $4(3)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for each $i\in[3]$ is a $3$-face.
$(a)$ Let $m_{4}(v)=0$. We first show that $v$ has at least one $8(7^{-})$-neighbours. Indeed, if all neighbours of $v$ are $7^{-}$- or 8(8)888(8)8 ( 8 )-vertices, then we would have $d_{2}(v)\leq 22$. When we set $G^{\prime}=G-v+\{v_{1}v_{4}\}$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction. Thus $v$ has at least one $8(7^{-})$-neighbours. First suppose that $v$ has exactly one $8(7^{-})$-neighbour. If $v$ has no $7(6^{-})$-neighbours, then we have $d_{2}(v)\leq 22$. Moreover, if $v$ has one $7(6^{-})$-neighbour, and two neighbours consisting of $6^{-}$-, 7(7)777(7)7 ( 7 )- or 8(8)888(8)8 ( 8 )-vertices, then we would have $d_{2}(v)\leq 22$. Furthermore, if $v$ has two $7(6^{-})$-neighbours and one neighbour consisting of $6^{-}$-, 7(7)777(7)7 ( 7 )- or 8(8)888(8)8 ( 8 )-vertices, then we would have $d_{2}(v)\leq 22$. Similarly as above, we get a contradiction. We then conclude that $v$ has three $7(6^{-})$-neighbours. Now, we suppose that $v$ has exactly two $8(7^{-})$-neighbours. If $v$ has no $7(6^{-})$-neighbour, then $v$ is adjacent to two 8(8)888(8)8 ( 8 )-neighbours, since otherwise, we have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. Thus, $v$ has either one $7(6^{-})$-neighbours or two 8(8)888(8)8 ( 8 )-neighbours.
$(b)$ Let $m_{4}(v)=1$. If $v$ has at most one $8(7^{-})$-neighbours, then we would have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. Thus, $v$ has at least two $8(7^{-})$-neighbours. Moreover, if $v$ has two neighbours consisting of $6^{-}$-, 7(7)777(7)7 ( 7 )- or 8(8)888(8)8 ( 8 )-vertices, then we would have $d_{2}(v)\leq 22$. Furthermore, if $v$ has two $8(7^{-})$-neighbours, one $7(6^{-})$-neighbour, and one neighbour consisting of $6^{-}$-, 7(7)777(7)7 ( 7 )- or 8(8)888(8)8 ( 8 )-vertices, then we have again $d_{2}(v)\leq 22$. Similarly as above, we get a contradiction. ∎
@@
Let $v$ be a 4(4)444(4)4 ( 4 )-vertex. Then, $v$ has either three $8(7^{-})$- and one $7(5^{-})$-neighbours or four $8(7^{-})$-neighbours.
@@
If $v$ has a $6^{-}$-, $7(6^{+})$- or 8(8)888(8)8 ( 8 )-neighbours, then we would have $d_{2}(v)\leq 22$. In addition, if $v$ has two $7^{-}$-neighbours, then we would have again $d_{2}(v)\leq 22$. When we set $G^{\prime}=G-v$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction. This implies that $v$ has either three $8(7^{-})$- and one $7(5^{-})$-neighbours or four $8(7^{-})$-neighbours. ∎
Lemma 2.34.
Let $v$ be a $5$-vertex.
  • $(a)$
    If $v$ is a 5(5)555(5)5 ( 5 )-vertex, and adjacent to a 5(5)555(5)5 ( 5 )-vertex, then $v$ has no $4$- and $5(4)$-neighbours.
  • $(b)$
    If $v$ is a 5(5)555(5)5 ( 5 )-vertex, then $v$ has at most one neighbour consisting of $4$- or 5(5)555(5)5 ( 5 )-vertices.
Proof.
$(a)$ Suppose that $v$ is a 5(5)555(5)5 ( 5 )-vertex adjacent to a 5(5)555(5)5 ( 5 )-vertex. If $v$ is also adjacent to a $4$- or $5(4)$-vertices, then we have $d_{2}(v)\leq 22$. When we set $G^{\prime}=G-v$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $v$ be a 5(5)555(5)5 ( 5 )-vertex. Suppose that $v$ has two neighbours consisting of $4$- or 5(5)555(5)5 ( 5 )-vertices. If $v$ has a 5(5)555(5)5 ( 5 )-neighbour, then we get a contradiction by $(a)$. Else, if $v$ has two $4$-neighbours, then we have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. ∎
@@
    If $m_{4}(v)=0$, then $v$ has two $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex.
@@
    If $m_{4}(v)=1$, then $v$ has three $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex.
@@
Let $v$ be a $5(4)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{5}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$.
$(a)$ Let $m_{4}(v)=0$. By contradiction, assume that $v$ has at most one $6^{+}$-neighbour different from 6(6)666(6)6 ( 6 )-vertex, i.e., all neighbours of $v$ but one consist of 6(6)666(6)6 ( 6 )- and $5^{-}$-vertices. Notice that if $v$ has a 6(6)666(6)6 ( 6 )-neighbour, then there exist two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces. It can be easily observed that $v$ has no four 6(6)666(6)6 ( 6 )-neighbours. Consider all the other possibilities; $v$ has either three 6(6)666(6)6 ( 6 )-neighbours or two 6(6)666(6)6 ( 6 )-neighbour and one $5^{-}$-neighbours or one 6(6)666(6)6 ( 6 )-neighbour and two $5^{-}$-neighbours or four $5^{-}$-neighbours. In each case, we have $d_{2}(v)\leq 22$. When we set $G^{\prime}=G-v+\{v_{1}v_{5}\}$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $m_{4}(v)=1$. Assume that $v$ has three neighbours $x,y,z$ consisting of $5^{-}$- or 6(6)666(6)6 ( 6 )-vertices. Note that if $v$ has a 6(6)666(6)6 ( 6 )-neighbour, then there exist two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces. Consider all the possibilities; all $x,y,z$ are $5^{-}$-vertices; two of $x,y,z$ are $5^{-}$-vertices and the last is 6(6)666(6)6 ( 6 )-vertex; one of $x,y,z$ is a $5^{-}$-vertex and the others are 6(6)666(6)6 ( 6 )-vertices; all $x,y,z$ are 6(6)666(6)6 ( 6 )-vertices. In each case, we have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. ∎
@@
    If $n_{4}(v)=1$, then $v$ has three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex.
@@
    If $n_{4}(v)=n_{5}(v)=0$, then $v$ has either three $8(7^{-})$-neighbours or four $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex such that two of which are $7^{+}$-vertices different from 7(7)777(7)7 ( 7 )-vertex.
@@
    If $n_{4}(v)=0$ and $n_{5}(v)=1$, then $v$ has either three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex or two $6(5^{-})$- and two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex.
@@
    If $n_{4}(v)=0$ and $n_{5}(v)=2$, then $v$ has three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex.
@@
Let $v$ be a 5(5)555(5)5 ( 5 )-vertex, and denote by $f_{1},f_{2},\ldots,f_{5}$ the $3$-faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in a cyclic fashion.
$(a)$ Let $n_{4}(v)=1$, and assume that $v$ has at most two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex, i.e., $v$ has two neighbours consisting of $5$-, $6$- or 7(7)777(7)7 ( 7 )-vertices. It is clear that $d_{2}(v)\leq 22$. When we set $G^{\prime}=G-v$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $n_{4}(v)=n_{5}(v)=0$. Assume that $v$ has at most two $8(7^{-})$-neighbours. If $v$ has at most three $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex, then $v$ has two 6(6)666(6)6 ( 6 )-neighbours as $n_{4}(v)=n_{5}(v)=0$. In particular, $v$ has a neighbour forming $7$- or 8(8)888(8)8 ( 8 )-vertices. In each case, we have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. Thus, $v$ has four $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex. If only one of those neighbours is a $7^{+}$-vertex different from 7(7)777(7)7 ( 7 )-vertex, then we again have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction.
$(c)$ Let $n_{4}(v)=0$ and $n_{5}(v)=1$. If $v$ has at most one $7^{+}$-neighbour different from 7(7)777(7)7 ( 7 )-vertex, then we have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. Thus, $v$ has at least two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex. Suppose that $v$ has exactly two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex. If $v$ has at most one $6(5^{-})$-neighbour, i.e., $v$ has one $5$-neighbour, at most one $6(5^{-})$ and one neighbour consisting of 6(6)666(6)6 ( 6 )- or 7(7)777(7)7 ( 7 )-vertices, then we have again $d_{2}(v)\leq 22$. Similarly as above, we get a contradiction.
$(d)$ Let $n_{4}(v)=0$ and $n_{5}(v)=2$. If $v$ has at most two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex, then we have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. ∎
@@
Let $v$ be a $6(5)$-vertex. Then, $v$ has at most four $5(4^{+})$-neighbours. In particular, if $m_{4}(v)=1$ and $v$ has four $5(4^{+})$-neighbours, then $v$ has two $8(6^{-})$-neighbours.
@@
Let $v$ be a $6(5)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{6}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$.
Assume for a contradiction that $v$ has five $5(4^{+})$-neighbours. This implies that there exist at least three edges in $G[N(v)]$ such that each of them is contained in two $3$-faces, and so we have $d_{2}(v)\leq 20$. If we set $G^{\prime}=G-v$, the graph $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $23$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction.
Now we suppose that $m_{4}(v)=1$, and $v$ has four $5(4^{+})$-neighbours. Since $v$ has four $5(4^{+})$-neighbours, there exist at least two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces. If $v$ has at most one $8(6^{-})$-neighbour, then we would have $d_{2}(v)\leq 22$. Similarly as above, we get a contradiction. ∎
Let $v$ be a $7^{+}$-vertex and let $x$ be a $5^{-}$-neighbour of $v$. If $v$ and $x$ have two common neighbours, i.e., the edge $vx$ is contained in two $3$-faces, then $x$ is called a support vertex of $v$.
@@
Let $v$ be a $7^{+}$-vertex. If $v$ has a support vertex, then $d_{2}(v)\geq 23$.
@@
Let $v$ be a $7^{+}$-vertex. Suppose that $v_{1}$ is a support vertex of $v$. By definition, $v_{1}$ is a $5^{-}$-vertex and it is adjacent to both $v_{2}$ and $v_{7}$. By contradiction, assume that $d_{2}(v)\leq 22$. When we set $G^{\prime}=G-v+\{v_{1}v_{3},v_{1}v_{4},v_{1}v_{5},v_{1}v_{6}\}$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction. ∎
We next apply discharging to show that $G$ does not exist. We use the same initial charges as before, with the following discharging rules.
@@
    Every $4(1)$-vertex receives $\frac{1}{6}$ from each of its $7(6^{-})$- and $8(7^{-})$-neighbours.
@@
    Let $v$ be a $4(2)$-vertex. Then,
@@
      $v$ receives $\frac{1}{6}$ from each of its $7(6^{-})$-neighbours,
@@
      $v$ receives $\frac{1}{4}$ from each of its $8(7^{-})$-neighbours.
  • R6:
    Let $v$ be a $4(3)$-vertex. Then,
@@
      $v$ receives $\frac{1}{6}$ from each of its $7(6^{-})$-neighbours,
@@
      $v$ receives $\frac{1}{3}$ from each of its $8(7^{-})$-neighbours,
@@
      $v$ receives $\frac{1}{9}$ from each of its 8(8)888(8)8 ( 8 )-neighbours.
  • R7:
    Every 4(4)444(4)4 ( 4 )-vertex receives $\frac{1}{3}$ from each of its $7(5^{-})$- and $8(7^{-})$-neighbours.
  • R8:
    Every $5(4)$-vertex receives $\frac{1}{9}$ from each of its $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex.
  • R9:
    Let $v$ be a 5(5)555(5)5 ( 5 )-vertex. Then,
@@
      $v$ receives $\frac{1}{9}$ from each of its $6(5^{-})$- and 7(7)777(7)7 ( 7 )-neighbours,
@@
      $v$ receives $\frac{2}{9}$ from each of its $7(6^{-})$- and $8$-neighbours.
  • R10:
@@
First we show that $\mu^{*}(f)\geq 0$ for each $f\in F(G)$. Given a face $f\in F(G)$, if $f$ is a $3$-face, then it receives $\frac{1}{3}$ from each of its incident vertices by R1, and so $\mu^{*}(f)\geq\ell(f)-4+3\times\frac{1}{3}=0$. If $f$ is a $4$-face, then $\mu(f)=\mu^{*}(f)=0$. Let $f$ be a $5^{+}$-face. By applying R2, $f$ sends $\frac{1}{5}$ to each of its incident vertices. It then follows that $\mu^{*}(f)\geq\ell(f)-4-5\times\frac{1}{5}\geq 0$. Consequently, $\mu^{*}(f)\geq 0$ for each $f\in F(G)$.

### v3

@@ 2.3. The case $\Delta=8$
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[3]$ be $3$-faces incident to $v$. Notice first that if $d_{2}(v)\leq 22$, then $G^{\prime}=G-v+\{v_{1}v_{4}\}$ a graph that is proper with respect to $G$, contradicting Remark 2.1. Thus we assume that $d_{2}(v)\geq 23$.
$(a)$ Let $m_{4}(v)=0$. We first show that $v$ has at least one $8(7^{-})$-neighbour. Indeed, if all neighbours of $v$ were $7^{-}$- or $8(8)$-vertices, then $d_{2}(v)\leq 22$ by Lemma 2.2, a contradiction. Thus $v$ has at least one $8(7^{-})$-neighbour.
Suppose first that $v$ has exactly one $8(7^{-})$-neighbour. If $v$ has no $7(6^{-})$-neighbours, then $d_{2}(v)\leq 22$. If it has exactly one $7(6^{-})$-neighbour and two neighbours consisting of $6^{-}$-, $7(7)$-, or $8(8)$-vertices, then again $d_{2}(v)\leq 22$. Similarly, if $v$ has two $7(6^{-})$-neighbours and one neighbour consisting of $6^{-}$-, $7(7)$-, or $8(8)$-vertices, we again obtain $d_{2}(v)\leq 22$. Each case contradicts Remark 2.1. Hence when $v$ has exactly one $8(7^{-})$-neighbour, it must have three $7(6^{-})$-neighbours.
Suppose next that $v$ has exactly two $8(7^{-})$-neighbours. If $v$ has no $7(6^{-})$-neighbours, then $v$ must have two $8(8)$-neighbours; otherwise $d_{2}(v)\leq 22$. Thus, when $v$ has two $8(7^{-})$-neighbours, it has either one $7(6^{-})$-neighbour or two $8(8)$-neighbours.
$(b)$ Let $m_{4}(v)=1$. If $v$ had at most one $8(7^{-})$-neighbour, then $d_{2}(v)\leq 22$ by Lemma 2.2, yielding a contradiction. Thus $v$ has at least two $8(7^{-})$-neighbours. If $v$ had two neighbours consisting of $6^{-}$-, $7(7)$-, or $8(8)$-vertices, then again $d_{2}(v)\leq 22$. Similarly, if $v$ had two $8(7^{-})$-neighbours, one $7(6^{-})$-neighbour, and one neighbour consisting of $6^{-}$-, $7(7)$-, or $8(8)$-vertices, then $d_{2}(v)\leq 22$. Both situations contradict with our assumption. ∎
@@
Let $v$ be a $4(4)$-vertex. Then $v$ has either three $8(7^{-})$- and one $7(5^{-})$-neighbour, or four $8(7^{-})$-neighbours.
@@
Note that if $d_{2}(v)\leq 22$, then $G^{\prime}=G-v$ is proper with respect to $G$, contradicting Remark 2.1. Thus we assume that $d_{2}(v)\geq 23$. If $v$ has a neighbour consisting of $6^{-}$-, $7(6^{+})$-, or $8(8)$-vertex, then $d_{2}(v)\leq 22$ by Lemma 2.2, a contradiction. Furthermore, if $v$ has two $7^{-}$-neighbours, then $d_{2}(v)\leq 22$ once again. Thus all neighbours of $v$ must be $8(7^{-})$- or $7(5^{-})$-vertices, and at most one neighbour can be of type $7(5^{-})$. Hence $v$ has either three $8(7^{-})$- and one $7(5^{-})$-neighbour, or four $8(7^{-})$-neighbours. ∎
@@
    If $m_{4}(v)=0$, then $v$ has two $6^{+}$-neighbours that are not $6(6)$-vertices.
@@
    If $m_{4}(v)=1$, then $v$ has three $6^{+}$-neighbours that are not $6(6)$-vertices.
@@
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$.
$(a)$ Assume $m_{4}(v)=0$, and suppose that $v$ has at most one $6^{+}$-neighbour other than a $6(6)$-vertex. Thus all but one neighbour of $v$ are $6(6)$- or $5^{-}$-vertices. Observe that $v$ cannot have four $6(6)$-neighbours. We now check all possibilities: three $6(6)$-neighbours; two $6(6)$- and one $5^{-}$-neighbour; one $6(6)$- and two $5^{-}$-neighbours; or four $5^{-}$-neighbours. In every case $d_{2}(v)\leq 22$ by Lemma 2.2. Setting $G^{\prime}=G-v+\{v_{1}v_{5}\}$ yields a graph that is proper with respect to $G$, contradicting Remark 2.1. Therefore $v$ must have at least two $6^{+}$-neighbours distinct from $6(6)$-vertices.
$(b)$ Assume $m_{4}(v)=1$, and suppose that $v$ has three neighbours $x,y,z$ that are either $5^{-}$- or $6(6)$-vertices. Checking the possibilities (all of $x,y,z$ are $5^{-}$-vertices; two $5^{-}$- and one $6(6)$-vertices; one $5^{-}$- and two $6(6)$-vertices; or all $6(6)$-vertices) gives $d_{2}(v)\leq 22$ in every case by Lemma 2.2. This again contradicts Remark 2.1. Thus $v$ must have at least three $6^{+}$-neighbours that are not $6(6)$-vertices. ∎
@@
    If $v$ is adjacent to another $5(5)$-vertex, then $v$ has no $4$- or $5(4)$-neighbours.
@@
    $v$ has at most one neighbour that is a $4$- or $5(5)$-vertex.
Proof.
$(a)$ Suppose that $v$ is a $5(5)$-vertex adjacent to a $5(5)$-vertex. If $v$ also has a $4$- or $5(4)$-neighbour, then $d_{2}(v)\leq 21$ by Lemma 2.2. Setting $G^{\prime}=G-v$ yields a graph that is proper with respect to $G$, contradicting Remark 2.1. Thus $v$ has no such neighbours.
$(b)$ Suppose that $v$ has two neighbours consisting of $4$- or $5(5)$-vertices. If one of them is a $5(5)$-vertex, part (a) already gives a contradiction. Otherwise, $v$ has two $4$-neighbours, implying $d_{2}(v)\leq 22$ by Lemma 2.2. As above, this yields a contradiction. Thus $v$ has at most one such neighbour. ∎
Lemma 2.38.
Let $v$ be a $5(5)$-vertex.
  • $(a)$
    If $n_{4}(v)=1$, then $v$ has three $7^{+}$-neighbours that are not $7(7)$-vertices.
  • $(b)$
    If $n_{4}(v)=n_{5}(v)=0$, then $v$ has either three $8(7^{-})$-neighbours or four $6^{+}$-neighbours different from $6(6)$-vertices such that two of which are $7^{+}$-vertices different from $7(7)$-vertex.
@@
    If $n_{4}(v)=0$ and $n_{5}(v)=1$, then $v$ has either three $7^{+}$-neighbours different from $7(7)$-vertex or two $6(5^{-})$- and two $7^{+}$-neighbours different from $7(7)$-vertex.
@@
    If $n_{4}(v)=0$ and $n_{5}(v)=2$, then $v$ has three $7^{+}$-neighbours different from $7(7)$-vertex.
@@
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in cyclic order.
$(a)$ Assume $n_{4}(v)=1$ and that $v$ has at most two $7^{+}$-neighbours different from $7(7)$-vertices. Then $v$ has two neighbours consisting of $5$-, 66-, or $7(7)$-vertices, which forces $d_{2}(v)\leq 22$ by Lemma 2.2. Removing $v$ yields a graph that is proper with respect to $G$, contradicting Remark 2.1. Thus $v$ must have three such $7^{+}$-neighbours.
$(b)$ Let $n_{4}(v)=n_{5}(v)=0$. Assume that $v$ has at most two $8(7^{-})$-neighbours. If $v$ has at most three $6^{+}$-neighbours other than $6(6)$-vertices, then it must have two $6(6)$-neighbours. This forces $d_{2}(v)\leq 22$ by Lemma 2.2, a contradiction. Hence $v$ has four $6^{+}$-neighbours different from $6(6)$-vertices. Among these four, if only one is a $7^{+}$-vertex different from $7(7)$-vertex, then again $d_{2}(v)\leq 22$ by Lemma 2.2, yielding the same contradiction. Therefore at least two must be such $7^{+}$-vertices.
$(c)$ Let $n_{4}(v)=0$ and $n_{5}(v)=1$. If $v$ has at most one $7^{+}$-neighbour different from $7(7)$-vertex, then $d_{2}(v)\leq 22$, a contradiction. Thus $v$ has at least two such $7^{+}$-neighbours.
Assume now that $v$ has exactly two $7^{+}$-neighbours different from $7(7)$-vertex. If $v$ has at most one $6(5^{-})$-neighbour, then again $d_{2}(v)\leq 22$ by Lemma 2.2. Thus $v$ must have two $6(5^{-})$-neighbours in this configuration.
$(d)$ Let $n_{4}(v)=0$ and $n_{5}(v)=2$. If $v$ has at most two $7^{+}$-neighbours different from $7(7)$-vertices, then $d_{2}(v)\leq 22$, again contradicting Remark 2.1. Hence $v$ has three such neighbours. ∎
@@
Let $v$ be a $6(5)$-vertex. Then $v$ has at most four $5(4^{+})$-neighbours. In particular, if $m_{4}(v)=1$ and $v$ has four $5(4^{+})$-neighbours, then $v$ has two $8(6^{-})$-neighbours.
@@
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$. Assume first that $v$ has five $5(4^{+})$-neighbours. Then at least three edges of $E(v)$ are contained in two $3$-faces, and Lemma 2.2 gives $d_{2}(v)\leq 20$. Since $v$ has five $5$-neighbours, one of $v_{1},v_{6}$ must be a $5$-vertex, say $v_{1}$, and one of $v_{3},v_{5}$ is also a $5$-vertex, say $v_{3}$. Let $G^{\prime}=G-v+\{\,v_{1}v_{6},\ v_{3}v_{1},\ v_{3}v_{5}\,\}$. Clearly, $G^{\prime}$ is proper with respect to $G$, contradicting Remark 2.1. Thus, $v$ has at most four $5(4^{+})$-neighbours.
Now suppose that $m_{4}(v)=1$ and that $v$ has four $5(4^{+})$-neighbours. Then there are at least two edges in $E(v)$ each contained in two $3$-faces. If $v$ has at most one $8(6^{-})$-neighbour, then $d_{2}(v)\leq 22$ by Lemma 2.2, and the same reduction as above yields a contradiction. Hence $v$ must have at least two $8(6^{-})$-neighbours. ∎
Let $v$ be a $7$-vertex (resp. an $8$-vertex), and let $x$ be a $5^{-}$-neighbour (resp. a $4^{-}$-neighbour) of $v$. If $v$ and $x$ have two common neighbours, (equivalently, if the edge $vx$ lies in two $3$-faces), then $x$ is called a support neighbour of $v$.
@@
Let $v$ be a $7^{+}$-vertex.
  • $(a)$
    If $v$ has a support neighbour, then $d_{2}(v)\geq 23$.
  • $(b)$
    If $v$ is a $7(6)$-vertex, and has five $4(1^{+})$-neighbours, then $v$ cannot have any $5$-neighbour.
  • $(c)$
    If $v$ is a $7(6)$-vertex, and has a $5(5)$-neighbour, then $v$ has at most four neighbours consisting of $4$- or $5(4)$-vertices.
  • $(d)$
    If $v$ is a $7(6)$-vertex, and has two $5(5)$-neighbours, then $v$ has at most three neighbours consisting of $4$- or $5(4)$-vertices.
  • $(e)$
    If $v$ is a $7(6)$-vertex, and has three $5(5)$-neighbours, then $v$ has at most one $4$-neighbour. In particular, $v$ has at most two neighbours consisting of $4$- or $5(4)$-vertices.
  • $(f)$
    If $v$ is a $7(7)$-vertex, then $v$ has at most five $5(4^{+})$-neighbours.
  • $(g)$
    If $v$ is a $8(7)$-vertex, and has a support neighbour, then $v$ has at most six neighbours consisting of $4(2^{-})$- or $5(5)$-vertices.
@@
$(a)$ Let $d(v)=k$ and suppose that $v_{1}$ is a support neighbour of $v$. By definition, $v_{1}$ is adjacent to both $v_{2}$ and $v_{k}$. Assume for a contradiction that $d_{2}(v)\leq 22$. Consider $G^{\prime}=G-v+\{v_{1}v_{3},v_{1}v_{4},\ldots,v_{1}v_{k-1}\}.$ The graph $G^{\prime}$ is proper with respect to $G$, contradicting Remark 2.1. Thus, $d_{2}(v)\geq 23$.
$(b)$ Suppose that $v$ is a $7(6)$-vertex, and $v$ has five $4(1^{+})$-neighbours. Clearly, $v$ has a support neighbour. If $v$ had a $5$-neighbour, then it would be $d_{2}(v)\leq 21$, contradicting (a).
$(c)$ Suppose that $v$ is a $7(6)$-vertex, and $v$ has a $5(5)$-neighbour. Clearly, $v$ has a support neighbour. If $v$ had five neighbours consisting of $4$- or $5(4)$-vertices, then it would be $d_{2}(v)\leq 22$, contradicting (a).
$(d)$ Suppose that $v$ is a $7(6)$-vertex with two $5(5)$-neighbours. Clearly, $v$ has a support neighbour. If $v$ had four neighbours consisting of $4$- or $5(4)$-vertices, then $d_{2}(v)\leq 22$, contradicting part (a).
$(e)$ Suppose that $v$ is a $7(6)$-vertex with three $5(5)$-neighbours. Clearly, $v$ has a support neighbour. Since a $5(5)$-vertex is adjacent to at most one $5(5)$-vertex by Lemma 2.37(b), $v$ can have at most one $4$-neighbour; otherwise $d_{2}(v)\leq 22$, contradicting (a). Similarly, $v$ cannot have three neighbours from $\{4\text{-vertices},\,5(4)\text{-vertices}\}$, as this would again imply $d_{2}(v)\leq 22$, contradicting (a).
$(f)$ Assume, for a contradiction, that $v$ has six $5(4^{+})$-neighbours. Clearly, at least one of such $5(4^{+})$-neighbour is a support neighbour of $v$, and Lemma 2.2 then gives $d_{2}(v)\leq 21$, contradicting (a).
$(g)$ Suppose that $v$ has seven such neighbours. Then Lemma 2.2 yields $d_{2}(v)\leq 22$, again contradicting (a). ∎
We now apply discharging to show that $G$ does not exist. We use the same initial charges as before, together with the following discharging rules.
@@
    Every $4(4)$-vertex receives $\frac{1}{3}$ from each of its $7(5^{-})$-neighbours.
@@
  • R6:
    Every $6(5^{-})$-vertex gives $\frac{1}{9}$ to each of its bad $5$-neighbours.
  • R7:
    Let $v$ be a $7(6^{-})$-vertex. Then, $v$ gives
@@
      $\frac{1}{6}$ to each of its $4(1)$-, $4(2)$- and $4(3)$-neighbours,
@@
      $\frac{1}{9}$ to each of its $5(4)$-neighbours,
    • $(c)$
      $\frac{2}{9}$ to each of its $5(5)$-neighbours,
  • R8:
    Every $7(7)$-vertex gives $\frac{1}{9}$ to each of its bad $5$-neighbours,
  • R9:
    Let $v$ be a $8(7^{-})$-vertex. Then, $v$ gives
@@
      $\frac{1}{6}$ to each of its $4(1)$-neighbours,
@@
      $\frac{1}{4}$ to each of its $4(2)$-neighbours,
@@
      $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours,
    • $(d)$
      $\frac{1}{9}$ to each of its $5(4)$-neighbours,
    • $(e)$
      $\frac{2}{9}$ to each of its $5(5)$-neighbours,
  • R10:
    Let $v$ be an $8(8)$-vertex. Then, $v$ gives
@@
      $\frac{1}{9}$ to each of its $4(3)$-, and 5(4)5(4)–neighbours,
@@
      $\frac{2}{9}$ to each of its $5(5)$-neighbours,
@@
First we show that $\mu^{*}(f)\geq 0$ for each $f\in F(G)$. Given a face $f\in F(G)$, if $f$ is a $3$-face, then it receives $\frac{1}{3}$ from each of its incident vertices by R1, and so $\mu^{*}(f)=\ell(f)-4+3\times\frac{1}{3}=0$. If $f$ is a $4$-face, then $\mu(f)=\mu^{*}(f)=0$. Let $f$ be a $5^{+}$-face. By applying R2, $f$ sends $\frac{1}{5}$ to each of its incident vertices. It then follows that $\mu^{*}(f)\geq\frac{4\ell(f)}{5}-4\geq 0$. Consequently, $\mu^{*}(f)\geq 0$ for each $f\in F(G)$.